Suppose a 1.5 μm wide slit produces its first minimum for 410 nm violet light.

Randomized Variables
λv = 410 nm
λr = 730 nm
w = 1.5 μm


(a) Calculate the angle at which this occurs for violet light in degrees.
(b) Where is the first minimum (in degrees) for 730 nm red light?
1
Expert's answer
2020-11-30T14:57:14-0500

(a) The angle at which this occurs for violet light θ = angular separation

For the 1st minimum n = 1

"sinθ = \frac{λ_v}{w}"

"θ = sin^{-1}(\frac{λ_v}{w})"

"θ = sin^{-1}(\frac{410 \times 10^{-9}}{1.5 \times 10^{-6}}) \\

= sin^{-1}(0.2733) \\

= 15.86 \; degrees"

(b) For the red light:

"dsinθ = nλ_r \\

n = 1 \\

sinθ = \frac{λ_r}{w} \\

θ = sin^{-1}(\frac{λ_r}{w}) \\

θ = sin^{-1}(\frac{730 \times 10^{-9}}{1.5 \times 10^{-6}}) \\

= sin^{-1}(0.4866) \\

= 29.11 \; degrees"